OpenStudy (anonymous):

The volume of a square pyramid varies jointly with its height and the square of its base side length. A square pyramid with a height of 9 meters and base side length 5 meters has a volume of 75 cubic meters. What is the base side length in meters of a square pyramid whose height is 12 meters and whose volume is 100 cubic meters?

OpenStudy (anonymous):

Dude you shouldn't have posted what you got...

OpenStudy (anonymous):

x = 18.75

OpenStudy (dumbcow):

no your work looked good side length = 5

OpenStudy (anonymous):

That has to be wrong

OpenStudy (anonymous):

How can the side length of both be right...I object

OpenStudy (anonymous):

nope told you lol

OpenStudy (anonymous):

dumb cow, how about you show your work....

OpenStudy (dumbcow):

they are proportional same base, diff heights yield diff volumes

OpenStudy (anonymous):

exactly as i told you before heromiles

OpenStudy (anonymous):

I don't agree

OpenStudy (anonymous):

dumbcow, show your work

OpenStudy (dumbcow):

V = 1/3*h*s^2 100 = (1/3)*12*s^2 100 = 4s^2 25 = s^2 5 = s

OpenStudy (anonymous):

and my work is here

OpenStudy (anonymous):

That's the same thing the other guy did, but I don't think so

OpenStudy (anonymous):

it is fine could someone help me with my question now please

OpenStudy (anonymous):

dude your wrong

OpenStudy (dumbcow):

google volume of a pyramid v=1/3*AreaBase*height

OpenStudy (anonymous):

I already know what the volume of a pyramid is...the question you have to ask yourself is, did you do it right...

OpenStudy (dumbcow):

i think so, if not please show me

OpenStudy (anonymous):

dumbcow come help me plz

OpenStudy (anonymous):

he is too stubborn to admit he is wrong lol

OpenStudy (anonymous):

I didn't know I was wrong silver...I verified my answer with what I though was a legitmate source..

OpenStudy (dumbcow):

prove it to yourself, think about it 18.5 ^2 = 342 which is much bigger than volume of 100

OpenStudy (anonymous):

yep, I admit I was wrong, you happy Silver?

OpenStudy (anonymous):

no, im just glad you learned something today